# Boundary of the Moduli Space of Stable Cubic Fivefolds

### v6

@@ Abstract
We study the boundary of the GIT quotient for cubic fivefolds from three complementary viewpoints: irreducible components, normal forms of closed orbits, and singularities. We give explicit normal forms representing closed (polystable) $\mathrm{SL}(7)$-orbits that appear on the boundary, and describe the geometry of the singular locus of each normal form (including the appearance of low-degree curves).
A key feature of the strictly semistable locus is that isolated singularities occur on the GIT boundary: we show that there are exactly six analytic types of isolated singularities appearing on strictly semistable cubic fivefolds, each given by a quasi-homogeneous normal-form family (some with moduli). For each of these six types we compute the local minimal exponent and prove that it is always equal to $\frac{7}{3}$. This value is precisely the critical threshold $\frac{n+1}{d}$ for degree $d=3$ hypersurfaces in $\mathbb{P}^{n}$ with $n=6$. In particular, these six isolated singularities realize the equality (critical) case in Park’s criterion for minimal exponents. Motivated by this, we call these six types extremal cubic fivefold singularities.
We further obtain strong computational evidence for the non-inclusion among the $21$ boundary families. Assuming Conjecture 3.5 (supported by these computations), it follows that the boundary decomposes into $21$ irreducible components. Finally, using wall-crossing in Kirwan’s stratification, we determine the adjacency relations among the boundary families; in particular, we show that the only nontrivial adjacencies occur in eight pairs.
@@ Introduction
Irreducible components and normal forms. We first construct explicit polystable (closed-orbit) normal forms and the corresponding boundary families $\Phi_{k}$ ($1\leq k\leq 21$), together with their parameter spaces and dimensions. Assuming Conjecture 3.5 (which is strongly supported by extensive numerical experiments; see Remark 3.6), these families give exactly the $21$ irreducible components of the strictly semistable locus; see Theorem A. In any case, the construction provides a concrete coordinate-level model for the boundary representatives and their degenerations.
Singular loci on the boundary. Next, we determine the singular locus of each closed-orbit representative $X_{k}$. We show that $\mathrm{Sing}(X_{k})$ consists of a one-dimensional part (when present) together with finitely many isolated singular points. Moreover, the positive-dimensional part is a union of low-degree curves—namely lines, smooth conics, complete intersections of two quadrics, or an elliptic quartic—whereas isolated singularities occur only in finitely many families. Crucially, all isolated boundary singularities are quasi-homogeneous and fall into precisely six analytic types. These statements are summarized in Theorem C.
@@
Adjacency of boundary families. Finally, we analyze adjacency relations among the boundary families arising from the closed-orbit normal forms. In a setting with many boundary families, the “connectivity” and “proximity” of families control the local structure of the moduli space and are closely tied to wall-crossing phenomena. Using Kirwan’s stratification and a detailed analysis of degenerations among the closed-orbit representatives, we show that the nontrivial adjacencies are extremely sparse: there are exactly eight adjacent pairs, as stated in Theorem D. This provides a minimal combinatorial skeleton for the boundary, and serves as a foundation for further investigations of local deformation models and related compactifications.
Idea of the proofs. Our approach combines standard GIT techniques—$1$-parameter-subgroup limits, computations of stabilizers and centralizers, and normalizations via residual group actions—with explicit calculations tailored to cubic fivefolds. For each boundary family we construct concrete normal forms and read off the structure of the singular locus directly from these models. The minimal exponent statement is obtained by applying Park’s criterion to our explicit list of isolated singularities and verifying the critical equality $\widetilde{\alpha}_{x}(X)=\frac{7}{3}$. The adjacency results follow from comparing degenerations along Kirwan strata and tracking how the boundary normal forms specialize.
@@
together with explicit closed-orbit (polystable) representatives in normal form; see Table 2 (Section 4). Assuming Conjecture 3.5 (strongly supported by extensive numerical experiments; see Remark 3.6), these families are pairwise distinct and give exactly the $21$ irreducible components of $\mathbb{P}(W)^{ss}\setminus\mathbb{P}(W)^{s}$.
@@
Theorem C (Singularities on the boundary). Let $\phi^{\mathrm{nf}}_{k}$ be the closed–orbit representative associated with $\Phi_{k}$ (Table 2) and set $X_{k}:=V(\phi^{\mathrm{nf}}_{k})\subset\mathbb{P}^{6}$. Then $\mathrm{Sing}(X_{k})$ is explicitly determined (Table 3): it is a union of low–degree curves (lines, smooth conics, complete intersections of two quadrics, or an elliptic quartic), possibly together with isolated points. Isolated points occur exactly for $k\in\{1,5,6,8,9,11,16,19,21\}$; they are quasi–homogeneous and fall into six analytic types (Definitions 5.1–5.6), including wild examples, and satisfy the extremal equality
@@
Theorem D (Adjacency via wall-crossing). Let $\overline{\Phi}_{k}\subset\mathbb{P}(W)//\mathrm{SL}(7)$ denote the image of the boundary family $\Phi_{k}$, and write $\Sigma^{\mathrm{ss}}:=\mathbb{P}(W)^{\mathrm{ss}}\setminus\mathbb{P}(W)^{\mathrm{s}}$ for the strictly semistable locus. We record adjacencies of boundary components in the sense of Kirwan’s partial desingularization, i.e. via wall-crossings between Kirwan strata. Then the adjacency graph on $\{\overline{\Phi}_{k}\}_{k=1}^{21}$ has exactly eight edges, namely
$\{\overline{\Phi}_{1},\overline{\Phi}_{7}\},\ \{\overline{\Phi}_{2},\overline{\Phi}_{6}\},\ \{\overline{\Phi}_{3},\overline{\Phi}_{12}\},\ \{\overline{\Phi}_{8},\overline{\Phi}_{19}\},\ \{\overline{\Phi}_{9},\overline{\Phi}_{15}\},\ \{\overline{\Phi}_{10},\overline{\Phi}_{17}\},\ \{\overline{\Phi}_{11},\overline{\Phi}_{21}\},\ \{\overline{\Phi}_{14},\overline{\Phi}_{20}\}.$
In particular, the remaining five components
$\overline{\Phi}_{4},\ \overline{\Phi}_{5},\ \overline{\Phi}_{13},\ \overline{\Phi}_{16},\ \overline{\Phi}_{18}$
are isolated vertices (i.e. adjacent to no other component). Equivalently, for $i\neq j$ one has
$\overline{\Phi}_{i}\cap\overline{\Phi}_{j}\cap\Sigma^{\mathrm{ss}}\neq\varnothing$
if and only if $\{i,j\}$ is one of the eight unordered pairs listed above. See Theorem 6.1 and [Kir84, DH98, Tha96].
@@
Our starting point is a convex-geometric analysis of Hilbert–Mumford weights for the natural $\mathrm{SL}(7)$-action on $W=\mathrm{Sym}^{3}(\mathbb{C}^{7})$ [MFK94, Dol03]. Fix a maximal torus $T\subset\mathrm{SL}(7)$, and for a cubic $f\in W$ write $\mathrm{Supp}(f)\subset I=\mathbb{Z}^{7}_{(3)}\cap\mathbb{Z}^{7}_{\geq 0}$ for its set of exponent vectors. We enumerate the maximal $T$-strictly semistable supports $I(r)_{\geq 0}$ whose convex hull contains the barycenter $\eta=(3/7,\dots,3/7)$. An algorithmic search over the faces of $\mathrm{Conv}(I)$ containing $\eta$ produces $23$ candidates up to permutation; exactly one is $T$-unstable and is discarded, leaving $22$ families with respect to $T$. Passing to the full $\mathrm{SL}(7)$-action, there is a single residual identification $f_{21}\sim f_{22}$. Extensive computations strongly suggest that no further identifications occur; see Conjecture 3.5 and Remark 3.6. This leads to the desired list of 2121 $\mathrm{SL}(7)$-inequivalent boundary families (Algorithm 2.2, Proposition 3.4, and Theorem 3.7).
@@
Section 1 recalls the numerical criterion for (semi)stability in convex-geometric language. Section 2 enumerates maximal strictly semistable supports for the maximal torus $T$ (Algorithm 2.2). Section 3 passes from $T$-data to the $\mathrm{SL}(7)$-action, eliminates redundancies, and arrives at the $21$ families $f_{1},\dots,f_{21}$ Section 4 constructs closed-orbit representatives, proves polystability (via the convex-hull or Casimiro–Florentino criterion), and records component dimensions. Section 5 computes singular loci and lists isolated types, including wild examples. Section 6 records adjacencies as wall-crossings in Kirwan’s stratification. Table 2 presents a compact summary of the normal forms and dimensions.
@@
This article is a preliminary arXiv version intended to quickly disseminate the main results. The scripts used to produce the computational data (including Tables 2–4) will be made publicly available in the next version of this preprint. In the meantime, the corresponding outputs are fully recorded in the paper.
@@
The author used ChatGPT and Gemini as auxiliary tools to assist in exploring and organizing computations, including the derivation of candidate normal forms in Section 4 and candidate normal forms of isolated singularities in Section 5. They were also used to support coding related to these computations (e.g., drafting and debugging scripts). In addition, these tools were used to improve exposition, organization, and internal consistency of the manuscript through iterative discussions. Under the author’s supervision, they were used for language editing in English. All mathematical statements and results were ultimately checked and verified by explicit calculations carried out by the author, who takes full responsibility for the content.
@@ Acknowledgments
The author would like to thank Professor T. Terasoma for his encouragement, advice, and valuable suggestions. He is deeply grateful to his parents for their support during difficult times. The author dedicates this work to the memory of Professor E. Horikawa.
@@ 2 Maximal strictly semistable cubic fivefolds with respect to the maximal torus $T$
Let $\mathcal{F}$ be the set of five different points of $I$. We fix a total order on $\mathcal{F}$. As an initial data, we set $\mathcal{S}^{\prime}=\varnothing$ and $\mathbf{x}=(x_{0},\cdots,x_{5})$ be the minimum element of $\mathcal{F}$. We will modify $\mathcal{S}^{\prime}$ using the following algorithm.
  • Step 1. If the subspace $W$ spanned by $x_{0},\cdots,x_{5},\eta$ of $\mathbb{Q}^{7}$ has dimension $6$ then take a reduced normal vector $\mathbf{r}=(r_{0},\cdots,r_{6})\in\mathbb{Z}^{7}_{(0)}$ of $W$ and go to Step 2, else go to Step 5.
  • Step 2. If $r=(r_{0},\cdots,r_{6})$ satisfies the condition $r_{0}\geq\cdots\geq r_{6}$ or $r_{0}\leq\cdots\leq r_{6}$, then go to Step 3, else go to Step 5.
  • Step 3. If $r_{0}\geq\cdots\geq r_{6}$ (resp. $r_{0}\leq\cdots\leq r_{6}$) add $I(\mathbf{r})$ (resp. $I(\mathbf{-r})$) to $\mathcal{S}^{\prime}$ and go to Step 4.
  • Step 4. Delete all elements of $\mathcal{S}^{\prime}$ that are not maximal in $\mathcal{S}^{\prime}$ and go to Step 5.
@@
Remark 3.6 (Computational evidence).
Extensive numerical experiments strongly suggest that, for any $k\neq l$, there is no $g\in\mathrm{SL}(7)$ satisfying
$g\cdot I(\mathbf{r}_{k})_{\geq 0}\subseteq I(\mathbf{r}_{l})_{\geq 0}.$
@@
Assuming Conjecture 3.5, the strictly semistable locus
@@
has 2121 irreducible components, represented by $f_{1},\dots,f_{21}$.

### v7

@@ Abstract
We give a description of the strictly semistable boundary of the GIT moduli space of cubic fivefolds. Specifically, we classify all $21$ boundary components and determine the codimension-one wall adjacency relation among them in the sense of Kirwan wall-crossing. The resulting adjacency graph is computed explicitly: it has $21$ vertices and $56$ edges (in particular, it has no isolated vertices). For each boundary component, we compute the minimal exponent of a general member and prove that it equals $7/3$, which is the critical value appearing in the stability criterion for cubic fivefolds.
Our approach is explicit: we exhibit $21$ normal forms representing the general polystable member in each boundary component and analyze the corresponding singularities. In particular, we identify exactly six analytic types of isolated singularities occurring on the boundary—extremal cubic fivefold singularities—each given by a quasi-homogeneous normal form (in some cases with moduli) and each having minimal exponent $7/3$. We also treat the non-isolated cases: we classify the possible positive-dimensional singular loci—lines, conics, complete intersections of two quadrics, and an elliptic quartic curve—and compute their minimal exponents.
@@ Introduction
Irreducible components and normal forms. We first construct explicit polystable (closed-orbit) normal forms and the corresponding boundary families $\Phi_{k}$ ($1\leq k\leq 21$), together with their parameter spaces and dimensions. We then prove the non-inclusion among these $21$ families (Theorem 3.5); in particular, the families $\Phi_{k}$ are pairwise distinct and give exactly the $21$ irreducible components of the strictly semistable locus; see Theorem A. In any case, the construction provides a concrete coordinate-level model for the boundary representatives and their degenerations.
Singular loci on the boundary. Next, we determine the singular locus of a general closed-orbit representative $X_{k}$ in each boundary family. We show that $\mathrm{Sing}(X_{k})$ consists of a one-dimensional part (when present) together with finitely many isolated singular points. Moreover, the positive-dimensional part is a union of low-degree curves—namely lines, smooth conics, complete intersections of two quadrics, or an elliptic quartic—whereas isolated singularities occur only in finitely many families. Crucially, on these general representatives, all isolated boundary singularities are quasi-homogeneous and fall into precisely six analytic types. These statements are summarized in Theorem C.
@@
Global minimal exponents on the boundary. The extremal value $\frac{7}{3}$ is not confined to isolated points. We compute minimal exponents along all positive-dimensional singular loci that occur on strictly semistable closed orbits (lines, conics, complete intersections of two quadrics, and an elliptic quartic), and show that possible drops occur only at finitely many explicitly described special points. As a consequence, the global minimal exponent of a general strictly semistable closed-orbit cubic fivefold in each boundary component is equal to the same critical threshold $7/3$ (see Theorem D).
Adjacency of boundary families. Finally, we analyze codimension-one wall adjacency relations among the boundary families arising from the closed-orbit normal forms. In a setting with many boundary families, their connectivity and proximity control the local structure of the moduli space and are closely tied to Kirwan-type wall-crossing. Using Kirwan’s stratification together with an explicit slice-maximality computation on the weight set $I=\mathbb{Z}^{7}_{\geq 0}(3)$, we determine the full codimension-one wall-adjacency graph among the closed strata $\{\Phi_{k}\}_{k=1}^{21}\subset\Sigma^{\mathrm{ss}}$. Contrary to a sparse picture, the resulting adjacency graph is nontrivial: it has $21$ vertices and $56$ edges (Theorem E and Theorem 7.2), and in particular it has no isolated vertices. For clarity, we distinguish genuine wall adjacencies (mutual slice matches, drawn as solid edges in Figure 2) from one-sided slice matches (drawn as dotted arrows), which do not represent codimension-one adjacency but rather reflect higher-codimension incidence among faces. Altogether, this provides an explicit combinatorial skeleton for the boundary, and serves as a foundation for further investigations of local deformation models and related compactifications.
Idea of the proofs. Our approach combines standard GIT techniques—$1$-parameter-subgroup limits, computations of stabilizers and centralizers, and normalizations via residual group actions—with explicit calculations tailored to cubic fivefolds. For each boundary family we construct concrete normal forms and read off the structure of the singular locus directly from these models. The minimal exponent statement is obtained by applying Park’s criterion to our explicit list of isolated singularities and verifying the critical equality $\widetilde{\alpha}_{x}(X)=\frac{7}{3}$. The adjacency results follow from comparing degenerations along Kirwan strata and tracking how the boundary normal forms specialize. Finally, the non-inclusion among the boundary families is established through a successive sieve of geometric and algebraic filters, utilizing the semicontinuity of dimensions, apolar Betti numbers, singular-locus Hilbert functions, minimal exponents, and Steenbrink spectra.
Organization. The paper provides a concrete “dictionary” for the GIT boundary of cubic fivefolds: the complete list of components, explicit closed-orbit normal forms, the associated singularity types, the critical minimal-exponent characterization, and the full adjacency graph. We hope that this dictionary will serve as a basis for subsequent work on the geometry of degenerations of cubic fivefolds, local deformation theory near boundary points, and comparisons with other notions of stability.
@@
together with explicit closed-orbit (polystable) representatives in normal form; see Table 2 (Section 4). Moreover, we prove the non-inclusion among these $21$ families (Theorem 3.5). In particular, the families $\Phi_{k}$ are pairwise distinct and give exactly the $21$ irreducible components of $\mathbb{P}(W)^{ss}\setminus\mathbb{P}(W)^{s}$.
@@
Theorem C (Singularities on the boundary). Let $\phi^{\mathrm{nf}}_{k}$ be a general closed–orbit representative associated with $\Phi_{k}$ (Table 2) and set $X_{k}:=V(\phi^{\mathrm{nf}}_{k})\subset\mathbb{P}^{6}$. Then $\mathrm{Sing}(X_{k})$ is explicitly determined (Table 3): it is a union of low–degree curves (lines, smooth conics, complete intersections of two quadrics, or an elliptic quartic), possibly together with isolated points. Isolated points occur exactly for $k\in\{1,5,7,8,9,11,16,19,21\}$; they are quasi–homogeneous and fall into six analytic types (Definitions 5.1–5.6), including wild examples, and satisfy the extremal equality
@@
Theorem D (Global minimal exponents on the boundary). For each $k\in\{1,\dots,21\}$ let $X_{k}:=V(\phi^{\mathrm{nf}}_{k})\subset\mathbb{P}^{6}$ be a general closed–orbit cubic fivefold. Then the global minimal exponent of $X_{k}$,
$\widetilde{\alpha}(X_{k})\ :=\ \min_{x\in\mathrm{Sing}(X_{k})}\widetilde{\alpha}_{x}(X_{k}),$
is equal to the extremal threshold value
$\widetilde{\alpha}(X_{k})=\frac{7}{3}\qquad\text{for all }k=1,\dots,21.$
More precisely, if $\mathrm{Sing}(X_{k})$ contains isolated points, then by Theorem C each such point $x$ satisfies $\widetilde{\alpha}_{x}(X_{k})=7/3$, hence $\widetilde{\alpha}(X_{k})=\frac{7}{3}$. If $\mathrm{Sing}(X_{k})$ contains positive–dimensional components (lines, conics, $(2,2)$-complete intersections, or an elliptic quartic), then $\widetilde{\alpha}_{x}(X_{k})$ is constant at a general point of each irreducible component, while it may drop at finitely many special points; in every case the minimum along each component is attained at an explicitly listed special point and equals $7/3$. Consequently, across the entire singular locus one has
$\min_{x\in\mathrm{Sing}(X_{k})}\widetilde{\alpha}_{x}(X_{k})=\frac{7}{3}.$
In particular, a general strictly semistable closed-orbit cubic fivefold in each of the 21 irreducible boundary components realizes the equality case
$\widetilde{\alpha}(X_{k})=\frac{n+1}{d}\qquad\text{for }(n,d)=(6,3).$
Theorem E (Codimension-one wall adjacency via wall-crossing). Write $\Sigma^{\mathrm{ss}}:=\mathbb{P}(W)^{\mathrm{ss}}\setminus\mathbb{P}(W)^{\mathrm{s}}$ for the strictly semistable locus. In the sense of Kirwan wall-crossing (Definition 7.1), the codimension-one wall-adjacency graph among the closed strata $\{\Phi_{k}\}_{k=1}^{21}\subset\Sigma^{\mathrm{ss}}$ has $21$ vertices and $56$ edges. Equivalently, for each $k$ the neighbor set
$N(k):=\{\,j\neq k\mid\Phi_{k}\ \text{is adjacent to}\ \Phi_{j}\,\}$
is given explicitly in Table 6 (and visualized in Figure 2). All other unordered pairs $\{\Phi_{i},\Phi_{j}\}$ are not adjacent. In particular, the adjacency graph has no isolated vertices. See Theorem 7.2 and [Kir84, DH98, Tha96].
@@
Our starting point is a convex-geometric analysis of Hilbert–Mumford weights for the natural $\mathrm{SL}(7)$-action on $W=\mathrm{Sym}^{3}(\mathbb{C}^{7})$ [MFK94, Dol03]. Fix a maximal torus $T\subset\mathrm{SL}(7)$, and for a cubic $f\in W$ write $\mathrm{Supp}(f)\subset I=\mathbb{Z}^{7}_{(3)}\cap\mathbb{Z}^{7}_{\geq 0}$ for its set of exponent vectors. We enumerate the maximal $T$-strictly semistable supports $I(r)_{\geq 0}$ whose convex hull contains the barycenter $\eta=(3/7,\dots,3/7)$. An algorithmic search over the faces of $\mathrm{Conv}(I)$ containing $\eta$ produces $23$ candidates up to permutation; exactly one is $T$-unstable and is discarded, leaving $22$ families with respect to $T$. Passing to the full $\mathrm{SL}(7)$-action, there is a single residual identification $f_{21}\sim f_{22}$. We prove that no further identifications occur: namely, we establish the non-inclusion among the $21$ boundary families (Theorem 3.5; see Section 8). This leads to the desired list of 2121 $\mathrm{SL}(7)$-inequivalent boundary families (Algorithm 2.2, Proposition 3.4, and Theorem 3.6).
@@
Section 1 recalls the numerical criterion for (semi)stability and reformulates it in convex–geometric terms via supports and weight half-spaces. Section 2 enumerates the maximal $T$-strictly semistable supports for a fixed maximal torus $T\subset\mathrm{SL}(7)$ (Algorithm 2.2), yielding $22$ candidates up to permutation. Section 3 passes from the $T$-data to the full $\mathrm{SL}(7)$-action, removes the single residual identification, and obtains 2121 $\mathrm{SL}(7)$-inequivalent boundary families $\Phi_{1},\dots,\Phi_{21}$ (with generic forms $f_{1},\dots,f_{21}$), together with the non-inclusion statement deferred to Section 8. Section 4 constructs explicit closed-orbit (polystable) representatives $\phi^{\mathrm{nf}}_{k}$ for each family via $1$-PS limits and Luna’s centralizer reduction, certifies polystability by either the convex-hull criterion (toric centralizers) or the Casimiro–Florentino criterion (non-toric centralizers), and computes the dimensions of the corresponding components (summarized in Table 2). Section 5 determines the singular scheme $\mathrm{Sing}(X_{k})$ of each closed-orbit cubic $X_{k}:=V(\phi^{\mathrm{nf}}_{k})\subset\mathbb{P}^{6}$, describing its positive-dimensional part (low-degree curves) and listing the isolated singularities; in particular, the isolated boundary singularities fall into exactly six quasi-homogeneous analytic types (summarized in Tables 3–4). Section 6 studies minimal exponents: it first proves that each of the six isolated boundary singularity types has local minimal exponent $7/3$, and then computes the global minimal exponent $\tilde{\alpha}(X_{k})$ for every $k=1,\dots,21$ by a case-by-case analysis along $\mathrm{Sing}(X_{k})$. Section 7 determines the adjacency relations among the closed strata $\{\Phi_{k}\}$ via wall-crossing in Kirwan’s stratification, showing that the adjacency graph has exactly 56 edges. Section 8 proves the non-inclusion among the $21$ boundary families (Theorem 3.5) by a sieve of monotone necessary conditions (“filters”), culminating in fine geometric and algebraic obstructions based on limits of $1$-cycles, Steenbrink spectra of isolated singularities, and generic stabilizer tori that remove the final candidates. Section 9 collects concluding remarks and outlines several future directions, including deformation/K-moduli questions for the extremal singularities and Hodge-theoretic and integrable-systems perspectives.
@@
The scripts used in this paper are publicly available at [Shi26]. In particular, the archive includes the scripts for Sections 2, 5, 7, 8.
@@
The author used ChatGPT and Gemini as auxiliary tools in several parts of the project. In Section 4, they were used to help explore and organize computations leading to candidate normal forms. In Section 5, they were used in the same way to help derive and organize candidate normal forms for isolated singularities. In Section 6, they were used to support the organization and exploration of the minimal exponent computations (e.g., structuring case-by-case checks and assisting with the interpretation and presentation of intermediate computational outputs). These tools were also used to support coding related to these computations (for example, drafting and debugging scripts). Finally, under the author’s supervision, they were used for English language editing and for improving exposition, organization, and internal consistency of the manuscript through iterative discussions. All mathematical statements and results were ultimately checked and verified by explicit calculations carried out by the author, who takes full responsibility for the content.
@@ Acknowledgments
The author is grateful to S. G. Park for valuable correspondence and insightful comments. The author would like to thank Professor T. Terasoma for his encouragement, advice, and valuable suggestions. He is deeply grateful to his parents for their support during difficult times. The author dedicates this work to the memory of Professor E. Horikawa.
@@ 2 Maximal strictly semistable cubic fivefolds with respect to the maximal torus $T$
Let $\mathcal{F}$ be the set of all ordered $5$-tuples $\mathbf{x}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{5})$ of pairwise distinct points of $I$. Fix a total order on $\mathcal{F}$. As initial data, set $\mathcal{S}^{\prime}=\varnothing$ and let $\mathbf{x}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{5})$ be the minimum element of $\mathcal{F}$. We modify $\mathcal{S}^{\prime}$ by the following algorithm.
  • Step 1. Let $W\subset\mathbb{Q}^{7}$ be the $\mathbb{Q}$-subspace spanned by $\mathbf{x}_{1},\ldots,\mathbf{x}_{5},\eta$. If $\dim_{\mathbb{Q}}(W)=6$, then take a reduced normal vector $\mathbf{r}=(r_{0},\ldots,r_{6})\in\mathbb{Z}^{7}_{(0)}\setminus\{0\}$ of $W$ (i.e. $\mathbf{r}\cdot\mathbf{w}=0$ for all $\mathbf{w}\in W$) and go to Step 2; otherwise, go to Step 5.
  • Step 2. If $\mathbf{r}$ satisfies $r_{0}\geq\cdots\geq r_{6}$ or $r_{0}\leq\cdots\leq r_{6}$, then go to Step 3; otherwise, go to Step 5.
  • Step 3. If $r_{0}\geq\cdots\geq r_{6}$ (resp. $r_{0}\leq\cdots\leq r_{6}$), add $I(\mathbf{r})_{\geq 0}$ (resp. $I(-\mathbf{r})_{\geq 0}$) to $\mathcal{S}^{\prime}$ and go to Step 4.
  • Step 4. Delete all elements of $\mathcal{S}^{\prime}$ that are not maximal in $\mathcal{S}^{\prime}$ (with respect to inclusion), and go to Step 5.
@@
Proof.
We postpone the proof of this theorem to Section 8. ∎
@@
The strictly semistable locus
@@
has $21$ irreducible components, represented by $f_{1},\dots,f_{21}$.